# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ Abstract
Optimization is one of the keystones of modern science and engineering. Its applications in quantum technology and machine learning helped nurture variational quantum algorithms and generative AI respectively. We propose a general approach to design variational optimization algorithms based on generative models: the Variational Generative Optimization Network (VGON). To demonstrate its broad applicability, we apply VGON to three quantum tasks: finding the best state in an entanglement-detection protocol, finding the ground state of a 1D quantum spin model with variational quantum circuits, and generating degenerate ground states of many-body quantum Hamiltonians. For the first task, VGON greatly reduces the optimization time compared to stochastic gradient descent while generating nearly optimal quantum states. For the second task, VGON alleviates the barren plateau problem in variational quantum circuits. For the final task, VGON can identify the degenerate ground state spaces after a single stage of training and generate a variety of states therein.
@@ I Introduction
Mathematical optimization is ubiquitous in modern science and technology. Spanning diverse fields like economics, chemistry, physics, and various engineering areas, its applications abound [31]. In quantum information theory, many problems relating to the approximation and characterization of quantum correlations can be formulated as convex optimization [4] problems, which is a particular kind of mathematical optimization with provable global optimality guarantees. Typical examples include the Doherty-Parrilo-Spedalieri (DPS) hierarchy [12] for entanglement certification and the Navascués-Pironio-Acín (NPA) hierarchy [45, 58] for bounding the set of quantum correlations, both relying on finding a hierarchy of semidefinite programming (SDP) relaxations of the original problems which converge to their optimal solutions.
@@
For quantum problems where convexity is hard to come by or the global optimality of the solution is a secondary consideration when compared to the efficiency of the algorithm, variational optimization provides a rich toolbox. In quantum many-body simulation, classical variational algorithms based on tensor network states allow the classical descriptions of the ground states of a large class of quantum models to be computed efficiently [9, 61, 16, 41, 64]. When solutions are expected to be quantum, hybrid quantum-classical variational algorithms are popular choices. Variational optimization is carried out on the classical parameters of these algorithms, while quantum gates and measurements are implemented in the corresponding quantum circuit [47, 38, 14, 55, 52, 20].
@@
Optimization is also at the core of every machine learning algorithm [43]. Generative models, especially the generative pretrained transformer (GPT) [48, 62] has opened a new way to address scientific problems that span a broad spectrum. In quantum physics, many-body quantum models can be efficiently solved by a generative model called the restricted Boltzmann machines [5, 40], lattice gauge theories can be simulated using another generative model called normalizing flows [34, 56], and quantum states can be more efficiently represented by variational autoencoders (VAEs) [29, 50, 53, 60, 35, 6].
@@
However, despite these encouraging advances, current applications of generative models to quantum problems usually focus on learning certain features from training data sets, and then generating new data with similar features. In a quantum problem, the training data may be a quantum state or complex correlation information contained therein, and the neural network is expected to generate new states or information resembling the training set. Extending the possibility of applying generative models to quantum problems beyond this scenario, we envisage using them to find (nearly) optimal solutions to general variational optimization problems.
Here we propose a general approach called the variational generative optimization network (VGON) to fulfill such a task. It contains a pair of classical deep feed-forward neural networks connected by a stochastic latent layer, and a problem-specific objective function. Two essential aspects of the VGON model include its integrated sampling mechanism within the training process and the adaptability of its objective function. The integrated sampling ensures that the outputs can be non-deterministic, which supports the identification of various simultaneous optimal solutions to the objective function. Meanwhile, the adaptable objective function enhances the model’s capacity to extend beyond simple replication of input data, indicating its wide applicability across diverse optimization challenges.
@@
We apply VGON to a variety of quantum problems to showcase its potential. We first demonstrate that it outperforms stochastic gradient descent (SGD) by avoiding entrapment in local optima in variational optimization problems of modest size, while also converging orders of magnitude faster. For larger problems with tens of thousands of parameters, we show that VGON can substantially alleviate the problem of barren plateaux in parameterized quantum circuits. Since generative models allow multiple optimal solutions to be found and generated simultaneously, a capability that deterministic algorithms lack, we use VGON to explore the ground state space of two quantum many-body models known to be degenerate. We show that VGON can successfully identify the dimension of the ground state space and generate a variety of orthogonal or linearly independent ground states spanning the entire space.
@@ II The VGON model
The architecture of VGON, shown in Fig. 1, is similar to that of the variational autoencoder. It consists of two deep feed-forward neural networks, the encoder $E_{\bm{\omega}}$ and the decoder $D_{\bm{\phi}}$, connected by a latent layer $\mathcal{Z}$ containing a normal distribution $\mathcal{N}(\bm{\mu(z)},{\bm{\sigma^{2}(z)}})$, where the mean $\bm{\mu}$ and the standard deviation $\bm{\sigma}$ are provided by $E_{\bm{\omega}}$. During the training stage, input data $\bm{x_{0}}$ is sampled from a distribution $P(\bm{x_{0}})$, which in all our tests is the uniform distribution over the parameter space. It is then mapped to the latent distribution $\mathcal{N}(\bm{\mu(z)},{\bm{\sigma^{2}(z)}})$ by the encoder network $E_{\bm{\omega}}$. Next the decoder network $D_{\bm{\phi}}(\bm{z})$ maps data $\bm{z}$ sampled from the latent distribution $\mathcal{N}(\bm{\mu(z)},{\bm{\sigma^{2}(z)}})$ to a distribution minimizing the objective function $h(\bm{x})$. This minimization is achieved by iteratively updating the parameters $\bm{\omega}$ and $\bm{\phi}$ in $E_{\bm{\omega}}$ and $D_{\bm{\phi}}$, respectively. Due to the existence of a stochastic latent layer, the gradients cannot be propagated backwards in the network. We solve this issue by using the reparameterization trick [29].
The key difference between VGON and VAE lies in the objective function (also called the loss function in the machine learning literature): instead of asking the output data distribution to approximate the input distribution by minimizing the Kullback-Leibler (KL) divergence between them, VGON simply requires the output data to minimize any objective function that is appropriate for the target problem. The decoder network, a versatile function approximator, maps the probability distribution from latent space to an output data distribution which can minimize the objective function with high probability [10, 11]. In addition, the KL divergence between the latent distribution $\mathcal{N}(\bm{\mu(z)},{\bm{\sigma^{2}(z)}})$ and the standard normal distribution $\mathcal{N}(\bm{0},\bm{I})$ is also minimized during training, as part of the objective function.
@@
Figure 1: The framework of VGON. The network is composed of an encoder network $E_{\bm{\omega}}$, a latent space $\mathcal{Z}$, and a decoder network $D_{\bm{\phi}}$. Training data $\bm{x_{0}}$ sampled from $P(\bm{x_{0}})$ is first mapped into a latent distribution $\mathcal{N}(\bm{\mu(z)},{\bm{\sigma^{2}(z)}})$ by $E_{\bm{\omega}}(\bm{x_{0}})$. Then a latent variable $\bm{z}$ sampled from $\mathcal{N}(\bm{\mu(z)},{\bm{\sigma^{2}(z)}})$ is transformed to the output $\bm{x}$ by $D_{\bm{\phi}}(\bm{z})$. The parameters $\bm{\phi}$ and $\bm{\omega}$ are updated iteratively to minimize the objective function $h(\bm{x})$.
@@
After the objective function converges to within a given tolerance, the training stage is completed. To utilize the trained model, the encoder network is disabled, and data sampled from a standard normal distribution $\mathcal{N}(\bm{0},\bm{I})$ are fed into the decoder network. Depending on the characteristics of the objective function, the corresponding output distribution can be tightly centered around one or multiple optimum values.
@@
To show that VGON can work well, we first use it to solve a variational optimization problem with a known unique optimal solution: finding the minimum ground state energy density among a class of quantum many-body models that matches the lower bound certified by an SDP relaxation. More specifically, we consider a class of infinite 1D translation-invariant (TI) models with fixed couplings discussed in Ref. [65], with the optimization variables being the local observables. The ground state energy density of this class of models has a lower bound certified by a variant of the NPA hierarchy [65]. However, there is no guarantee that any infinite TI quantum many-body Hamiltonian, if couplings are fixed but the dimension of local observables can be arbitrary, can achieve this bound. Meanwhile, by optimizing 3-dimensional local observables with SGD and computing the ground state with uniform matrix product state algorithms [41, 61], a Hamiltonian whose ground state energy density matches the above lower bound to 7 significant digits has been found. We replace SGD with VGON to conduct the same optimization, and find that the converged model can (almost) deterministically generate Hamiltonians whose ground state energy density matches the NPA lower bound to 8 significant digits, reaching the precision limit of commercial SDP solvers (see Appendix A for more technical details). Below we apply VGON to several more complicated problems.
@@ III Finding the optimal state for entanglement detection
In a prepare-and-measure entanglement detection game, what is the best state for experimentally demonstrating the advantage of local operations and one-way classical communication (1-LOCC) protocols over protocols using local operations (LO) alone? This problem has practical relevance and does not have a known optimal solution. It can be formulated as a variational optimization problem consisting of multiple convex optimization sub-problems, making it a good showcase of the power of VGON.
Suppose two players, Alice and Bob, receive a bipartite quantum state $\rho$ from a source and want to determine whether $\rho$ is entangled. They have a choice of either performing the experiment in their respective labs, then communicating the outcomes from Alice to Bob thus implementing a 1-LOCC protocol, or forgoing communication completely therefore making the protocol LO. These two scenarios have been formalized as preparation games and introduced in Refs. [63, 24] as a new way of automatically devising experimental protocols to detect quantum entanglement that can minimize the probability of committing statistical errors. Although on certain target quantum states it is possible to show 1-LOCC protocols lead to smaller error probabilities than any LO ones, the advantage is highly dependent on the choice of target states and is hard to analyze.
In this problem, the advantage is defined to be the difference between the minimum probabilities $p_{2}$ of committing false-negative statistical errors (a.k.a. type-II errors, defined as a source distributes an entangled state, but Alice and Bob conclude the state they received is separable). It can be computed by solving the following two SDPs and subtracting the two minimal $p_{2}^{*}$ corresponding to LO and 1-LOCC protocols:
$\min$ $p_{2}$ (1)
$s.t.$ ${\rm tr}(M_{N}\rho)=p_{2},\ p_{1}\mathbb{I}-M_{Y}\in\mathcal{S}^{*},$
$M_{Y}=M_{Y}(P),\ M_{N}=M_{N}(P),$
@@
Here $\mathcal{S}^{*}$ denotes the dual of the separable set $\mathcal{S}$. $p_{1}$ is the given probability of Alice and Bob making false-positive errors, i.e. type-I error, defined as the scenario in which they conclude that they have received an entangled state, while the source actually distributes a separable one. The positive operator-valued measure (POVM) operators $M_{Y(N)}$ can be constructed as $M_{Y(N)}=\sum_{x,y,a,b}P(x,y,Y(N)|a,b)\;A^{a}_{x}\otimes B^{b}_{y}$, where $\{A_{x}^{a}\}$ ($\{B_{y}^{b}\}$) are predetermined measurements performed by Alice (Bob) with $x$ ($y$) being measurement labels and $a$ ($b$) being outcomes. The choice of their protocols is reflected in the sets $\mathcal{P}^{LO}$ or $\mathcal{P}^{1-LOCC}$ in which $P(x,y,Y(N)|a,b)$ is constrained. A more detailed description of the theoretical framework can be found in Appendix B.
For a random quantum state $\rho$, it turns out that the gap calculated above is usually very small, as shown by the green dots in Figure 2(a). In order to observe the gap under noisy experimental conditions, we would like to know which $\rho$ exhibit the largest gap among the states that can be prepared in a general quantum linear optical setup. Using such a setup as a convenient way to parameterize the state space, we first employ SGD to optimize the gap by starting from random initial pure bipartite qutrit states. The results are shown in Figure 2(a). Because of being trapped easily in local maxima and of needing to compute the gradients with dozens of SDPs, it takes two months to conduct optimizations for 79,663 random initial pure states (see Appendix B.1 for more details). Most of these states exhibit gaps around $0.0036$ before optimization, while the largest gap afterwards is $0.083722$.
The results of using VGON to solve this problem are depicted in Figure 2(b) and summarized in Table 1. Based on 3,000 sets of initial parameters produced by uniform sampling, the model converges after less than two hours of training. After that, we use it to generate 100,000 output states. We find that over 98% of them manifest gaps over 0.08, while over 50% of them have gaps larger than $0.0835$. A similar performance has been observed even when choosing the initial quantum states from a variational submanifold of the space of all mixed states, where out of 100,000 states generated by the converged VGON model, 83 have gaps larger than $0.0837$, with an average purity of $0.99999$ (see Appendix B.2 for more details).
Figure 2: Comparing SGD and VGON for generating states with large gaps. (a) Most of the 79,663 random initial states for SGD exhibit small gaps around $0.0036$, while after optimization most states have gaps smaller than $0.08$. The largest gap is $0.083722$. (b) Over $98.59\%$ of the 100,000 states generated by a trained VGON model have gaps larger than $0.08$, with the largest group, consisting of over $50\%$ of the generated states, tightly centered around $0.0837$.
Table 1: Performance comparison between VGON and SGD on finding states maximizing the advantage of 1-LOCC protocols over LO ones.

### v2

@@ Abstract
Optimization drives advances in quantum science and machine learning, yet most generative models aim to mimic data rather than to discover optimal answers to challenging problems. Here we present a variational generative optimization network that learns to map simple random inputs into high quality solutions across a variety of quantum tasks. We demonstrate that the network rapidly identifies entangled states exhibiting an optimal advantage in entanglement detection when allowing classical communication, attains the ground state energy of an eighteen spin model without encountering the barren plateau phenomenon that hampers standard hybrid algorithms, and—after a single training run—outputs multiple orthogonal ground states of degenerate quantum models. Because the method is model agnostic, parallelizable and runs on current classical hardware, it can accelerate future variational optimization problems in quantum information, quantum computing and beyond.
@@ I Introduction
Mathematical optimization is ubiquitous in modern science and technology. Spanning diverse fields like economics, chemistry, physics, and various engineering areas, its applications abound [37]. In quantum information theory, many problems relating to the approximation and characterization of quantum correlations can be formulated as convex optimization problems [54, 71, 15], which is a particular kind of mathematical optimization with provable global optimality guarantees. For quantum problems where convexity is hard to come by or the global optimality of the solution is a secondary consideration when compared to the efficiency of the algorithm, variational optimization provides a rich toolbox. When solutions are expected to be quantum, hybrid quantum-classical variational algorithms are popular choices. In these algorithms, variational optimization is carried out on the classical parameters, while quantum gates and measurements are implemented in the corresponding quantum circuit [58, 45, 17, 68, 65, 24].
@@
Optimization is also at the core of every machine learning algorithm [51]. Recently, machine learning algorithms have opened a new way to address scientific problems spanning a broad spectrum, accelerating the integration of AI into the scientific discovery process [39, 76]. In mathematics, they help humans discover new results [11] and develop faster solutions to problems [18]. In biology, they help with drug developments [30]. Particularly, generative models have seen explosive growth in the form of large language models [60, 74], which are transforming the way humans interact with machines. Applying these models to science has enabled new solutions to mathematical problems to be discovered [63]. Meanwhile, generative models have also been widely applied to quantum physics. For example, many-body quantum models can be efficiently solved by restricted Boltzmann machines [7, 47], lattice gauge theories can be simulated using normalizing flows [41, 69], quantum states can be more efficiently represented by variational autoencoders (VAEs) [35, 62, 42, 8], and quantum circuits with desired properties can be generated by the generative pre-trained transformer [52].
@@
However, despite these encouraging advances, current applications of generative models to quantum problems usually focus on learning certain features from training data sets, and then generating new data with similar features. In the scenario where a classical (i.e., not quantum) generative model is used to solve a quantum problem, the training data may be quantum states or complex correlation information contained therein, and a neural network is expected to generate new quantum states or information resembling the training set.
In order to extend the possibility of applying generative models to quantum problems beyond this scenario, inspired by the classical variational autoencoder, we propose a method called the variational generative optimization network (VGON), whose output does not just resemble the input, but can be (nearly) optimal solutions to general variational optimization problems. VGON contains a pair of deep feed-forward neural networks connected by a stochastic latent layer, and a problem-specific objective function. The intrinsic randomness in the model can be leveraged both in its training and testing stages. During the training stage, we have not encountered any issues with the optimization getting trapped in local minima. We believe this can be partially explained by having random inputs, which effectively gives the optimization multiple starting points, and the architecture of our model, especially the existence of the latent layer, which regularizes the input and leads to good trainability. In the testing stage, the randomness allows VGON to produce multiple optimal solutions to the objective functions simultaneously, even after only a single stage of training.
@@
We apply VGON to a variety of quantum problems to showcase its potential. We first demonstrate that it outperforms stochastic gradient descent (SGD) by avoiding entrapment in local optima in variational optimization problems of modest size, while also converging orders of magnitude faster. For larger problems with tens of thousands of parameters, we show that VGON can substantially alleviate the problem of barren plateaux in parameterized quantum circuits. Since generative models allow multiple optimal solutions to be found and generated simultaneously, a capability that deterministic algorithms lack, we use VGON to explore the ground state space of two quantum many-body models known to be degenerate. We show that VGON can successfully identify the dimensionality of the ground state space and generate a variety of orthogonal or linearly independent ground states spanning the entire space.
@@ II The VGON model
The architecture of VGON, shown in Figure 1, consists of two deep feed-forward neural networks, the encoder $E_{\bm{\omega}}$ and the decoder $D_{\bm{\phi}}$ are connected via a latent layer $\mathcal{Z}$ containing a normal distribution $\mathcal{N}(\bm{\mu(z)},{\bm{\sigma^{2}(z)}})$, where the mean $\bm{\mu}$ and the standard deviation $\bm{\sigma}$ are provided by $E_{\bm{\omega}}$. During the training stage, input data $\bm{x_{0}}$ is sampled from a distribution $P(\bm{x_{0}})$, which in all our tests is the uniform distribution over the parameter space. It is then mapped to the latent distribution $\mathcal{N}(\bm{\mu(z)},{\bm{\sigma^{2}(z)}})$ by the encoder network $E_{\bm{\omega}}$. Next the decoder network $D_{\bm{\phi}}(\bm{z})$ maps data $\bm{z}$ sampled from the latent distribution $\mathcal{N}(\bm{\mu(z)},{\bm{\sigma^{2}(z)}})$ to a distribution minimizing the objective function $h(\bm{x})$. This minimization is achieved by iteratively updating the parameters $\bm{\omega}$ and $\bm{\phi}$ in $E_{\bm{\omega}}$ and $D_{\bm{\phi}}$, respectively. Due to the existence of a stochastic latent layer, the gradients cannot be propagated backwards in the network. We solve this issue by using the reparameterization trick [35].
@@
Figure 1: The framework of Variational Generative Optimization Network. The network is composed of an encoder network $E_{\bm{\omega}}$, a latent space $\mathcal{Z}$, and a decoder network $D_{\bm{\phi}}$. Training data $\bm{x_{0}}$ sampled from $P(\bm{x_{0}})$ is first mapped into a latent distribution $\mathcal{N}(\bm{\mu(z)},{\bm{\sigma^{2}(z)}})$ by $E_{\bm{\omega}}(\bm{x_{0}})$. Then a latent variable $\bm{z}$ sampled from $\mathcal{N}(\bm{\mu(z)},{\bm{\sigma^{2}(z)}})$ is transformed to the output $\bm{x}$ by $D_{\bm{\phi}}(\bm{z})$. The parameters $\bm{\phi}$ and $\bm{\omega}$ are updated iteratively to minimize the objective function $h(\bm{x})$, together with the Kullback-Leibler (KL) divervence between the latent distribution $\mathcal{N}(\bm{\mu(z)},{\bm{\sigma^{2}(z)}})$ and the standard normal distribution $\mathcal{N}(\bm{0},{\bm{I}})$.
The key difference between VGON and VAE lies in the objective function (also called the loss function in the machine learning literature): instead of asking the output data distribution to approximate the input distribution by maximizing a given similarity measure, VGON simply requires the output data to minimize any objective function that is appropriate for the target problem. In addition, the Kullback-Leibler (KL) divergence between the latent distribution $\mathcal{N}(\bm{\mu(z)},{\bm{\sigma^{2}(z)}})$ and the standard normal distribution $\mathcal{N}(\bm{0},\bm{I})$ is also minimized during training, as part of the objective function.
@@
After the objective function converges to within a given tolerance, the training stage is complete. To utilize the trained model, the encoder network is disabled, and data sampled from a standard normal distribution $\mathcal{N}(\bm{0},\bm{I})$ are fed into the decoder network. Depending on the characteristics of the objective function, the corresponding output distribution can be tightly centered around one or multiple optimum values. Moreover, it is worth mentioning that requiring the latent layer to follow a normal distribution not only facilitates efficient optimization of the objective function but also simplifies the sampling process, since the KL divergence between two normal distributions can be analytically evaluated and sampling from a normal distribution is computationally efficient.
The goal of VGON can be seen as finding a way which maps a simple distribution defined over the latent space, i.e., the Gaussian, to a complicated one, i.e., a distribution whose samples can minimize the objective function with high probability [14]. This shares the spirit of transport theory, where given two probability measures $\mu$ and $\nu$ on spaces $X$ and $Y$, we call a map $T\mathrel{\mathop{\ordinarycolon}}X\to Y$ a transport map if $T_{*}(\mu)=\nu$, where $T_{*}(\mu)$ is the pushforward of $\mu$ by $T$, representing the process of transferring (or “pushing forward”) the measure $\mu$ from $X$ to $Y$ via the measurable function $T$. In an optimal transport problem [75], one is interested in finding a map $T$ that minimizes the transport cost $\int c(x,T_{*}(x))d\mu(x)$, subject to the constraint that the pushforward measure satisfies $T_{*}(\mu)=\nu$ [59]. In addition to optimal transport problems, there are problems that do not have an explicitly defined target distribution, where the task is evaluating the loss function directly on the generated samples. In these problems, an optimal $T$ can either be found analytically, such as in inverse transform sampling, where both distributions are one-dimensional [12], or $T$ can be learned/optimized from a parameterized $T_{\theta}$ on training data, such as generative model like normalizing flows [41, 80], where both distributions are high-dimensional but $T$ is invertible and constructed using neural networks. In VGON, the latent space usually has a much lower dimension than the output layer, making $T$ surjective, which means every point in the target space can be reached by applying the decoder to some latent input [55]. This surjectivity relaxes the requirement for invertibility and enables VGON to easily cover the complex target distributions. In our experience, the best optimization results come from training the encoder-latent space-decoder triple as a whole, even though it is possible to achieve good results without including the encoder in the training process.
@@
To show that VGON can work well, we first use it to solve a variational optimization problem with a known unique optimal solution: finding the minimum ground state energy density among a class of quantum many-body models that matches the lower bound certified by an SDP relaxation [49, 72]. More specifically, we consider a class of infinite 1D translation-invariant (TI) models with fixed couplings [79], and the optimization variables are the local observables. The ground state energy density of this class of models has a lower bound certified by a variant of the NPA hierarchy [79]. However, there is no guarantee that any infinite TI quantum many-body Hamiltonian, if couplings are fixed but the local observables can be arbitrary, can achieve this bound. Meanwhile, by optimizing 3-dimensional local observables with SGD and computing the ground state with uniform matrix product state algorithms, a Hamiltonian whose ground state energy density matches the above lower bound to 7 significant digits has been found. We replace SGD with VGON to conduct the same optimization, and find that the converged model can (almost) deterministically generate Hamiltonians whose ground state energy density matches the NPA lower bound to 8 significant digits, reaching the precision limit of commercial SDP solvers (see Appendix Reaching the quantum limit of a many-body contextuality witness for more technical details). Below we apply VGON to several more complicated problems.
@@ III Finding the optimal state for entanglement detection
Entanglement detection plays a central role in quantum tasks such as secure communication and distributed computing, where entanglement serves as a fundamental resource. Suppose two players, Alice and Bob, receive a bipartite quantum state $\rho$ from a source, then they want to determine whether $\rho$ is entangled, with the smallest statistical error.
They can either perform the experiment independently in their respective laboratories and subsequently communicate the outcomes from Alice to Bob, or choose to forgo communication entirely. In the first scenario they are implementing a local operations and one-way classical communication (1-LOCC) protocol while in the second scenario they are implementing a local operations (LO) one. Implementing the 1-LOCC protocol experimentally requires fast real-time switching of Bob’s measurement settings and a quantum memory to store Bob’s half of the state while Alice performs her measurement and communicates the result. Do these extra experimental complexities yield tangible advantages such as reduced statistical error probabilities? In fact, it has been shown that for some simple states, such an advantage does exist, but it is too small to be useful [77]. In fact, the advantage is highly dependent on the choice of target states and is hard to estimate analytically. The goal of this task is to identify high-dimensional quantum states for which this advantage, defined as the gap between the minimum statistical error probabilities in LO and 1-LOCC protocols, is as large as possible. This gap quantifies the practical advantage of allowing one-way communication between the parties in entanglement detection.
Specifically, given a quantum state $\rho$, the advantage is defined to be the gap between the minimum probabilities $p_{2}$ of committing false-negative errors (a.k.a. type-II errors, defined as a source distributes an entangled state, but Alice and Bob conclude the state they received is separable) when Alice and Bob employ LO and 1-LOCC protocols, and the two protocols have the same probability of making false-positive errors, i.e., type-I error, denoted $p_{1}$ and defined as the scenario in which they conclude that they have received an entangled state, while the source actually distributes a separable one. The desired quantity can be computed by solving two SDPs that share the same objective but differ in constraint structure. Specifically, the following SDP is solved twice, once under LO protocol $P\in\mathcal{P}^{LO}$ and once under 1-LOCC protocol $P\in\mathcal{P}^{1\text{-}LOCC}$. The final result is obtained by subtracting the respective optimal values of $p_{2}$ [78]:
$\min_{P}$ $p_{2}$ (1)
$s.t.$ ${\rm tr}(M_{N}(P)\rho)=p_{2},$
$p_{1}\mathbb{I}-M_{Y}(P)\in\mathcal{S}^{*},$
@@
Here $\mathcal{S}^{*}$ denotes the dual of the separable set $\mathcal{S}$. The positive operator-valued measure (POVM) operators $M_{Y(N)}(P)$ can be constructed as $M_{Y(N)}(P)=\sum_{x,y,a,b}P(x,y,Y(N)|a,b)\;A^{a}_{x}\otimes B^{b}_{y}$, where $\{A_{x}^{a}\}$ ($\{B_{y}^{b}\}$) are predetermined measurements performed by Alice (Bob) with $x$ ($y$) being measurement labels and $a$ ($b$) being outcomes, and $P$ is the shorthand for the distribution $P(x,y,Y(N)|a,b)$, which specifies the detection strategy by assigning probabilities to particular combinations of settings, outcomes, and decisions. Here $Y$ and $N$ denote the decisions corresponding to the presence or absence of entanglement, respectively.
For a random quantum state $\rho$, it turns out that the gap calculated above is usually very small, as shown by the green dots in Figure 2(a). In order to observe the gap under noisy experimental conditions, We focus on a linear optical setup that generates bipartite qutrit states, which also allows us to parameterize the state space. We first employ SGD to maximize the gap by starting from random pure bipartite qutrit states. The results are shown in Figure 2(a). The SGD algorithm gets trapped easily in local maxima and needs to compute the gradient by solving dozens of SDPs. Optimizing the gap for 79,663 random pure states is computationally very costly (see Appendix The pure state case for more details). Most of these states exhibit gaps around $0.0036$ before optimization, while the largest gap afterwards is $0.083722$.
Figure 2: Comparisons of stochastic gradient descent (SGD) and Variational Generative Optimization Network (VGON) in generating states with large gaps. (a) Most of the 79,663 random initial states for SGD exhibit small gaps around $0.0036$, while after optimization $1.52\%$ of states have gaps larger than $0.08$, which is indicated by the dashed line. The largest gap is $0.083722$. (b) Over $98.59\%$ of the 100,000 states generated by a trained VGON model have gaps larger than $0.08$, which is presented by the dashed line. In particular, over $50\%$ of these states are tightly centered around $0.0837$.
Table 1: Performance comparison between Variational Generative Optimization Network (VGON) and stochastic gradient descent (SGD). The comparison is carried out on finding states which maximize the advantage of one-way local operations and classical communication (1-LOCC) protocols over local operations (LO) ones.
@@
The results of using VGON to maximize the gap are depicted in Figure 2(b) and summarized in Table 1. Based on 3,000 sets of initial parameters produced by uniform sampling, the model converges after less than two hours of training. After that, we use it to generate 10,000 output states. We find that over 98% of them manifest gaps over 0.08, while over 50% of them have gaps larger than $0.0835$. A similar performance has been observed even when choosing the initial quantum states from a variational submanifold of the space of all mixed states, where out of 10,000 states generated by the converged VGON model, 83 have gaps larger than $0.0837$, with an average purity of $0.99999$. For comparison, we also apply seven other global optimization algorithms and a Multilayer Perceptron (MLP)—a neural network with multiple fully connected layers—to this task, and find that VGON consistently outperforms all the baseline methods (see the second subsection in the Methods and Appendix Finding the optimal state for entanglement detection for more details). Importantly, using VGON to solve an experiment-relevant instance of this problem allows us to experimentally demonstrate the advantage of 1-LOCC protocols over LO ones in detecting entanglement, where we can observe an error probability that is impossible to achieve with local operations alone [78].